is log4 64^x = x true or false
Answer by stanbon(75887)

is log4 64^x = x true or false
------
log4(4^3)^x = x
---
log4(4^(3x)) = x
----
3x = x
----
True if x = 0
====================
Cheers,
Stan H.
=====================